# Talent Hoarding in Organizations

### v2

@@ Appendix E Supplementary Results for Robustness - For Online Publication
Table E2: Robustness in Effects of Manager Rotation on Worker Outcomes
@@
Notes: This table reports the effects of manager rotations on workers’ career outcomes with respect to applying and landing higher-level positions. In this table, higher-level positions are defined as those that have a hierarchy ranking of at least 10 based on methods developed in Haegele (2024). Each coefficient is based on a separate regression. Column 1 reports the first-stage effect of manager rotation on applications based on Equation 3. Column 2 reports estimates from a two-stages least squares regression on landing a position that instruments for applying with manager rotation based on Equation 5 which represents the LATE. Column 3 estimates a similar two-stages least squares regression, but use an indicator for landing a position and performing better than the leave-out team average one year later as outcome variable. Robust standard errors are clustered at the worker and rotation level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Figure E2: Mean Deviation Between Actual and Predicted Potential Ratings
Notes: This figure depicts the mean deviation by manager between predicted and actual potential ratings, which captures systematic discrepancies in worker visibility, and serves as a measure of managers’ propensities to hoard talent. The y-axis reports the number of unique managers with a given deviation. I assess managers’ systematic underreporting of public potential ratings by comparing managers’ actual potential rating to the predicted potential rating based on managers’ own assessment of worker performance and worker characteristics. Values above zero represent managers who on average lower their public signal below their private signal of worker talent, which captures one likely dimension of talent hoarding. I use the top (marked in orange) and bottom (marked in blue) tercile of this distribution to classify managers as high-propensity versus low-propensity to hoard talents. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects. N=6,xxx.
Table E3: Placebo Test for Talent Hoarding Measure
Panel A Panel B Panel C
Internal Applications Internal Transition External Transitions
High Low High Low High Low
(1) (2) (3) (4) (5) (6)
Manager Rotation 0.0344 0.0144 0.0166 0.0101 -0.0014 0.0020
(0.006) (0.005) (0.004) (0.003) (0.002) (0.002)
Outcome Mean 0.029 0.029 0.008 0.008 0.008 0.008
Observations 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx
Notes: This table provides a placebo test for my measure of managers’ propensities to hoard talent based on the mean deviation between actual and predicted potential ratings. Each coefficient stems from a separate regression based on Equation 3. Each panel compares rotations of managers with high versus low propensity to hoard. The outcomes of interest are internal applications (Panel A), internal job transitions within the firm (Panel B), and external job transitions out of the firm (Panel C). Robust standard errors are clustered at the worker and rotation level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
Table E4: Robustness in Hoarding and Quality Measure
Panel A: Measure of Managerial Talent Hoarding Propensity
Joint Sample Joint Sample Separate Sample
Bottom $\frac{1}{2}$ Top $\frac{1}{2}$ Bottom $\frac{1}{4}$ Top $\frac{1}{4}$ Bottom $\frac{1}{3}$ Top $\frac{1}{3}$
(1) (2) (3) (4) (5) (6)
Manager Rotation 0.0167 0.0287 0.0155 0.0320 0.0189 0.0323
(0.004) (0.004) (0.004) (0.005) (0.005) (0.006)
P-value of t-test 0.0771 0.0251 0.0084
Observations 3xx,xxx 3xx,xxx 3xx,xxx 3xx,xxx 1xx,xxx 1xx,xxx
Panel B: Measure of Worker Quality
Quality Index Education Performance
Bottom $\frac{1}{2}$ Top $\frac{1}{2}$ Bottom $\frac{1}{2}$ Top $\frac{1}{2}$ Bottom $\frac{1}{2}$ Top $\frac{1}{2}$
(1) (2) (3) (4) (5) (6)
Manager Rotation 0.0115 0.0349 0.0128 0.0277 0.0172 0.0278
(0.003) (0.005) (0.004) (0.004) (0.004) (0.004)
P-value of t-test 0.0000 0.0082 0.0628
Observations 1xx,xxx 1xx,xxx 1xx,xxx 1xx,xxx 1xx,xxx 2xx,xxx
Notes: This table documents the robustness of the heterogeneity analysis with respect to the application effects of manager rotations. Each coefficient stems from a separate regression based on Equation 3. Robust standard errors are clustered at the worker and rotation level. Panel A documents the robustness of the measure of talent hoarding based on deviations between managers’ actual and predicted potential ratings. Columns 1 and 2 split managers based on whether their hoarding measure is below or above the median instead of terciles used in my preferred measure. Columns 3 and 4 split managers based on the bottom and top quartiles of the hoarding measure. Columns 5 and 6 use terciles, but instead of classifying rotation events based on managers’ hoarding measure, I split workers into separate samples based on their managers’ hoarding measure. Panel B documents the robustness of my measure of worker quality. Columns 1 and 2 split workers into high-quality versus low-quality using the median of the quality index, instead of using quartiles as in my preferred approach. Columns 3 and 4 distinguish high-quality versus low-quality workers based on the median education level (having 18 or more years of schooling versus less than 18 years of schooling). Columns 5 and 6 make the quality distinction based on the median performance rating. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.

### v3

@@ E.3.4 Alternative Measures of Higher-Level Positions
Figure E3: Robustness of Rotation Effects to Alternative Measures of Worker Quality
Notes: This figure examines whether the effect of manager rotations on worker applications varies consistently across alternative measures of worker quality. The outcome is an indicator for whether a worker applied to an internal job in a given quarter. Each coefficient stems from a separate regression based on Equation 3. Panel A shows the baseline average effect of a manager rotation (in blue). The orange coefficients show the effect of the rotation interacted with a standardized (mean-zero, SD=1) version of the respective proxy for hoarding incentives. Panel B presents my baseline specification which uses a standardized composite index of worker quality (constructed as the predicted value from an OLS regression of internal hiring probabilities on worker characteristics). Panel C uses a standardized measure of years of schooling. Panel D uses a standardized measure of past performance ratings. Robust standard errors are clustered at the worker and rotation level. 95%-level confidence intervals are shown. Controls include: gender, age, German citizenship, education, marital and family status, parental leave, firm tenure, division, function, location, full-time status, hours worked, number of direct reports, and quarter fixed effects. N = 3xx,xxx.
Figure E4: Robustness to Alternative Measures of Manager Hoarding Propensity
Notes: This figure examines whether the effect of manager rotations on worker applications varies consistently across alternative specifications of the manager hoarding propensity measure. The outcome is an indicator for whether a worker applied to an internal job in a given quarter. Each coefficient stems from a separate regression based on Equation 3. Panel A shows the baseline average effect of a manager rotation (in blue). The orange coefficients show the interaction of the rotation effect with different versions of the manager hoarding propensity measure. Panel B uses the continuous version of the hoarding measure, defined as the average deviation between predicted and actual potential ratings. Panels C to E show results when discretizing this variable. Panel C uses an indicator for whether the manager is in the top tercile of the distribution, Panel D uses the top quartile, and Panel E uses the top half. In each case, the orange point represents the interaction of the rotation indicator with the respective indicator for being hoarding prone. Robust standard errors are clustered at the worker and rotation level. 95%-level confidence intervals are shown. Controls include: gender, age, German citizenship, education, marital and family status, parental leave, firm tenure, division, function, location, full-time status, hours worked, number of direct reports, and quarter fixed effects. N = 3xx,xxx.
Table E12: Robustness to Alternative Definition of Higher-Level Positions
@@
Notes: This table reports the effects of manager rotations on worker career outcomes using an alternative classification of higher-level positions based on job hierarchy rankings. Positions are classified as higher-level if they have a hierarchy rank of 10 or above, following the hierarchy measure developed in Haegele (2024), which incorporates number of direct reports, reporting distance to the CEO, and managerial autonomy. Each column presents results from a separate regression. Each coefficient is based on a separate regression. Column 1 reports the first-stage effect of manager rotation on applications based on Equation 3. Column 2 reports estimates from a two-stages least squares regression on landing a position that instruments for applying with manager rotation based on Equation 5 which represents the LATE. Column 3 estimates a similar two-stages least squares regression, but use an indicator for landing a position and performing better than the leave-out team average one year later as outcome variable. Robust standard errors are clustered at the worker and rotation level. Controls: Female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, division, function, location, full-time, hours, number of direct reports, and quarter fixed effects.
E.3.5 Alternative Mechanisms Underlying Rotation Design
Having demonstrated that the observed application effects of manager rotations are robust across a wide range of empirical specifications and measurement approaches, I now turn to an important question: could these effects be driven by mechanisms other than a reduction in talent hoarding?
The analysis in Section 5.2 is consistent with talent hoarding being a key driver of worker responses to manager rotations; however, it remains possible that worker responses could also reflect factors unrelated to talent hoarding. For instance, managers and workers may jump ship because of bad news about the future outlook of the team, or may depart after the completion of a major milestone. Other alternative mechanisms such as worker loyalty, match effects, and role-model effects could also produce the observed worker responses. In this section, I conduct a series of additional robustness checks to assess the potential importance of such alternative mechanisms.
Bad news about the future or unpleasant working conditions.— I begin by testing whether a negative shock, such as bad news about the future outlook of the unit or unpleasant working conditions, could underlie the effect of manager rotations on worker applications. Under such a mechanism, one would expect to see effects on other team-level outcomes besides applications, such as team performance and morale. Appendix Figure A2 estimates an event study around manager rotations and rejects economically significant changes in key outcomes, including team-level absenteeism and performance-related bonus pay, suggesting that rotations are not correlated with deteriorating conditions within the team.
Results from two placebo tests further contradict patterns that one would expect from such shocks. First, under a correlated shock that makes managers and workers want to leave the team, one would expect increased worker applications even for managers’ unsuccessful attempts to leave the team. However, Panel A of Appendix Figure E5 finds no effect of managers’ unsuccessful rotations on worker applications. Second, because coworkers are more similar in their career opportunities to a worker, one would also expect the correlation between coworker rotations and worker applications in the presence of such a shock to be larger than that between manager rotations and worker applications. Panel B of Appendix Figure E5 estimates the impacts of rotations of the most senior teammate and shows that these events have much smaller impacts on worker applications.
In addition, if the documented increase in internal applications is driven by teammates jumping ship, one would expect to see at least some effect on external transitions. However, I do not find that manager rotations yield increases in exit from the firm (Panel B of Figure 3). Moreover, if teammates want to jump ship because of the long-term future outlook of the team, one should expect the increase in applications to persist for several quarters, particularly given that in the internal labor market, it can take several quarters for suitable vacancies to become available. Instead, the observed application responses are transitory. It is also unclear why workers in smaller teams, those that are harder to replace, and those under hoarding-prone managers would be more susceptible to such a shock, as suggested by the heterogeneity results presented in Figure 4.353535Managers’ hoarding propensities are quite stable over time and are unlikely to be driven by a correlated shock.
Completion of a milestone or project.—If most teams were to work on fixed project schedules together, the completion of a milestone could be another key channel for correlated shocks. Several institutional features imply that such milestones are not likely to be relevant in this setting. First, manager rotations are not designed to coincide with increases in job vacancies or team dissolution that would indicate systematic rotation patterns due to common milestones in the firm. Second, the vast majority of positions at the firm are not project-based. Even in areas where project-based work is more common, projects are not normally staffed within a team, but often have employees from different teams in one project, meaning that it is not common that an entire team and their manager will finish a milestone at the same time. To empirically test for this channel, I exclude workers in jobs related to project management (where the bulk of project-based work occurs) from my event study analysis. This robustness test yields very similar application patterns (Panel B of Appendix Figure E6), suggesting that my results are not likely to be driven by project cycles. These findings are also consistent with the absence of effects on external transitions. If milestone completion were to drive responses to rotations, one would expect some effect on external transitions.
Loyalty.—In principle, manager rotations could induce workers to apply if they feel a strong sense of loyalty toward their manager. While loyalty undoubtedly represents an important part of interpersonal relations in the workplace, I do not find evidence that this mechanism drives the particular increase in applications around manager rotations. Loyalty is typically assumed to compound over time; however, Appendix Figure E7 finds no evidence that rotation effects vary by the length of time a worker was exposed to the rotating manager. Instead, rotations increase applications even for workers who have been exposed to their manager for one quarter or less. While one could have also expected that loyalty is particularly strong if the worker was personally hired by the manager, Appendix Table B5 finds similar effects for workers who were hired by the rotating manager and those who were already in the team. Moreover, one would expect loyalty effects to also result in increases in external transitions. However, Panel B of Figure 3 shows that while internal and external transitions trend identically prior to the rotation, manager rotations do not yield an increase in worker’s external transition out of the firm.
Worker-Manager Specific Match Effects.—Workers may have refrained from applying prior to a manager rotation because of match effects, which could occur because the manager provided favorable task assignments or offered better training opportunities. Several pieces of evidence suggest, however, that standard match effects cannot explain the increase in applications around a manager rotation. Similar to loyalty, match effects should increase worker applications for positions outside the team because the value of their default option falls when the manager leaves. However, I find no impact of manager rotations on external transitions. Under match effects, one would also expect rotations to be more impactful for workers who were hired by that manager, relative to workers who were already on the team when that manager arrived. This is because the manager would have had the ability to select new workers based on match quality. Instead, Appendix Table B5 documents that manager rotations have very similar effects on both groups. While one would also expect that match effects compound over time, I find that application effects of manager rotations do not depend on workers’ exposure time to their manager (Appendix Figure E7). Finally, it is difficult to rationalize why match effects would give rise to the observed heterogeneity across different workers, for instance in terms of larger rotation effects for workers under hoarding-prone managers. Evidence discussed in Section 4 suggests that match effects are likely smaller, not larger, under hoarding-prone managers, for example because they deter workers’ participation in trainings.
Salience and Role-Model Effects.—Information-based mechanisms could also explain why rotations affect worker applications. For instance, managers who pursue rotations may generate information flows that make career planning more salient. However, since managers’ unsuccessful rotation attempts do not increase applications even though workers likely find out about them, such a channel does not appear to operate through the salience of career planning. Successful manager rotations, on the other hand, could increase applications by providing information about the success likelihood of applications. If this were the case, one would expect that the increase in applications would persist, rather than quickly dissipate, given that few vacancies are available each quarter. Similarly, when I focus on manager rotations that are induced by the outgoing manager leaving the firm, which are less likely to increase information or salience about internal career transitions, I find similar application patterns relative to my baseline results (Panel A of Appendix Figure E6).
Role-model effects are often found to be particularly impactful if role models are similar in attributes to affected individuals (Porter and Serra, 2020, Meier et al., 2020), but homophily does not appear to play an important role in my setting. First, motivated by the previous literature on homophily by gender, I show that the rotation of a manager with the same gender yields a similar and statistically indistinguishable application response relative to the rotation of a manager with the opposite gender (Panel A of Appendix Table B7). Second, if observing others navigate their career is a key underlying factor of rotation effects, such a role-model effect should not be limited to managers. Since coworkers are more similar to workers than their managers with respect to their career opportunities, one would likely expect larger effects for observing coworkers rotate. However, Panel B of Appendix Figure E5 finds much smaller, rather than larger, effects of coworker rotations relative to manager rotations. These findings indicate that any role-model effects that are at play likely operate through channels that differ from those documented in prior research.
These additional analyses suggest that several plausible alternative explanations for the observed application effects around manager rotations are unlikely to be the dominant factors in my setting. Instead, the findings are in line with the view that reduced incentives to hoard talent are a primary driver of increased internal job applications.
Figure E5: Placebo Tests for Effects of Manager Rotations
Panel A. Results by Outcome of Manager Application
Panel B. Results by Type of Rotating Teammate
Figure E6: Effects of Alternative Rotation Events
Panel A.Manager Exits as Event
Panel B. Sample Without Project Management
Figure E7: Effect of Manager Rotations by Workers’ Exposure Length to Manager